Problem 2.1. Thinking about differentiating Taylor series, compute the sum n=0 for any z < 4.
Answer #1


Si Gwen senos Ž (nt. 90 n=0 = (0+1) Go? =1+2+7+36)* +4(70+ ---- =(1-4) 2 :(1+x)= 1+2x+3x74x + --.) = (4-772 15 (+x)= 1+2x+3